OpenStudy (anonymous):

HELP! file:///C:/Users/harry_000/Downloads/ggfuydh.PNG

OpenStudy (anonymous):

Copy the link and paste in a new tab. will FAN and METAL

OpenStudy (amistre64):

that is not how you attach a file ....

OpenStudy (anonymous):

How do you attach it?

OpenStudy (amistre64):

you can only attach a file after you open your post .. use the buttons below here vvvvvv

OpenStudy (anonymous):

click the button that says "attach file"

OpenStudy (anonymous):

also nice username c;

OpenStudy (amistre64):

much better, now people can help you out better :)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Do you have an answer?

OpenStudy (amistre64):

we are not an answering service

pooja195 (pooja195):

Let's start with #1 What step would you have to take to isolate k?

pooja195 (pooja195):

Isolate= get k alone

OpenStudy (anonymous):

What is k

pooja195 (pooja195):

We dont know and it doesnt matter we just want to get it alone

OpenStudy (anonymous):

Okay, what after that?

pooja195 (pooja195):

You tell me.

OpenStudy (alexandervonhumboldt2):

first do what pooja told you, then we will talk about what is next

pooja195 (pooja195):



pooja195 (pooja195):

How do we get rid of 6?

OpenStudy (anonymous):

-8.7=k

OpenStudy (anonymous):

Add 6 + -8.7????

OpenStudy (anonymous):

Or subtract

pooja195 (pooja195):

Coorect we subtract :) -8.7-6=k

OpenStudy (anonymous):

-14.7

pooja195 (pooja195):

We don't need to solve

OpenStudy (anonymous):

Okay

pooja195 (pooja195):

Which of your answer choices match?

OpenStudy (anonymous):

k = –8.7 – 6

pooja195 (pooja195):

Try #2 tell me which answer choice you think it is. Make sure you explain.

OpenStudy (anonymous):

I think –6 = k + 8.7 would be 8.7 – (–6) Because K is 8.7

OpenStudy (anonymous):

I mean k= is -6

pooja195 (pooja195):

Hmm try again... Flip the equation. k+8.7=−6 now subtract 8.7

pooja195 (pooja195):

What do you get?

OpenStudy (anonymous):

So,-6 - 8.7?

pooja195 (pooja195):

Correct

OpenStudy (anonymous):

-14.7

pooja195 (pooja195):

We don't need to know the values it only wants the 1st steps

pooja195 (pooja195):

Do the same for #3

OpenStudy (anonymous):

Just give me the answers

pooja195 (pooja195):

Nawh i'll pass